# Three Methods, One Problem: Classical and AI Approaches to No-Three-in-Line

### v1

@@ Abstract
The No-Three-In-Line problem asks for the maximum number of points that can be placed on an $n\times n$ grid with no three collinear, representing a famous problem in combinatorial geometry. While classical methods like Integer Linear Programming (ILP) guarantee optimal solutions, they face exponential scaling with grid size, and recent advances in machine learning offer promising alternatives for pattern-based approximation. This paper presents the first systematic comparison of classical optimization and AI approaches to this problem, evaluating their performance against traditional algorithms. We apply PatternBoost transformer learning and reinforcement learning (PPO) to this problem for the first time, comparing them against ILP. ILP achieves provably optimal solutions up to $19\times 19$ grids, while PatternBoost matches optimal performance up to $14\times 14$ grids with 96% test loss reduction. PPO achieves perfect solutions on $10\times 10$ grids but fails at $11\times 11$ grids, where constraint violations prevent valid configurations. These results demonstrate that classical optimization remains essential for exact solutions while AI methods offer competitive performance on smaller instances, with hybrid approaches presenting the most promising direction for scaling to larger problem sizes.
@@ 1 Introduction
A trivial upper bound of $2n$ points follows from the pigeonhole principle, as no row or column can contain more than two points. The first non-trivial lower bound was established by Paul Erdős, who used modular parabola constructions to place approximately $n$ points [ErdosRoth1951]. This was improved to approximately $1.5n$ by Hall, Jackson, Sudbery, and Wild using modular hyperbola constructions, which remains the best proven lower bound [Hall1975]. Exact solutions achieving $2n$ points have been found for all $n\leq 58$. However, Guy and Kelly conjectured that for sufficiently large $n$, the maximum asymptotically approaches only $cn$ points, where $c\approx 1.814$ [GuyKelly1968]. This leaves a significant gap between the best-proven result and the conjectured limit, underscoring the problem’s persistent difficulty.
@@ 2.4.2 PPO Algorithm and Network Architecture
The policy network uses a multi-layer perceptron (MLP) with two hidden layers of 512 units each and ReLU activations. The architecture processes the flattened $n^{2}$-dimensional observation vector to produce action probabilities over the $n^{2}$ action space. The policy network uses a multi-layer perceptron (MLP) with two hidden layers of 512 units each and ReLU activations. The architecture processes the flattened $n^{2}$-dimensional observation vector to produce action probabilities over the $n^{2}$ action space. The policy network (actor) $\pi_{\theta}(a|s)$ and value function (critic) $V_{\theta}(s)$ share feature extraction layers but have separate output heads, as shown in Figure 3.
@@ 3.2 Solution Quality and Scalability
Table 3 presents the solution quality achieved by each method across different grid sizes. We report the maximum number of points successfully placed while satisfying the no-three-in-line constraint (zero violations). For reference, we include known optimal values where available from the literature.
@@
Proximal Policy Optimization. PPO achieved perfect solutions on $n=10$ grids (20 points with zero violations), successfully learning valid placement policies through reward-based feedback. On $n=11$ grids, the agent placed 22 points but incurred a single collinearity violation, indicating difficulty maintaining long-term constraint satisfaction as problem complexity increases. PPO demonstrated the ability to discover effective strategies from minimal problem structure (only the reward signal), but struggled to scale beyond small grid sizes where sequential planning becomes critical. Figure 8 illustrates this sharp performance transition between successful constraint satisfaction at $n=10$ and systematic failure at $n=11$, with detailed analysis of failure modes provided in Section 3.4.
@@ 3.4 Limitations
Figure 8 illustrates PPO’s performance transition between successful and failed constraint satisfaction. On $10\times 10$ grids (Figure 8), the agent consistently achieves perfect solutions with 20 points and zero violations, demonstrating successful policy learning for problems of this scale. However, on $11\times 11$ grids (Figure 8), despite placing 22 points, the agent incurs a single collinearity violation (highlighted point in red box), revealing the brittleness of learned policies as problem complexity increases. The violation occurs in a densely populated region where multiple collinearity constraints intersect, suggesting that the policy network struggles to maintain global constraint awareness when the number of placed points exceeds approximately 20. This represents a sharp transition from reliable performance to systematic failure as grid size increases by just one row and column.
@@
Figure 8: PPO performance comparison showing success at $n=10$ and failure at $n=11$.
@@ 4 Discussion and Conclusion
Future research should prioritize hybrid training paradigms that combine supervised pretraining (PatternBoost’s cross-entropy learning) with reinforcement fine-tuning (PPO’s exploratory optimization), mirroring successful approaches in language modeling. Transfer learning across grid sizes and curriculum strategies could enable training on smaller grids before scaling to larger instances, with meta-learning potentially discovering optimization strategies that generalize across problem scales. Classical-AI integration offers additional promising directions: ILP warm-starting with PatternBoost candidates could accelerate convergence on larger grids, while PatternBoost training on ILP-optimal smaller solutions could improve pattern quality beyond greedy baselines. Physics-informed approaches present a particularly compelling direction, reformulating the No-Three-In-Line problem as energy minimization where Universal Differential Equations (UDEs) combine physics-based gradients with learned neural components, or Physics-Informed Neural Networks (PINNs) enforce geometric constraints through differentiable loss functions.

### v2

@@ Abstract
The No-Three-In-Line problem asks for the maximum number of points that can be placed on an $n\times n$ grid with no three collinear, representing a famous problem in combinatorial geometry. While classical methods like Integer Linear Programming (ILP) guarantee optimal solutions, they face exponential scaling with grid size, and recent advances in machine learning offer promising alternatives for pattern-based approximation. This paper presents a systematic comparison of classical optimization and AI approaches to this problem, evaluating their performance against traditional algorithms. We apply PatternBoost transformer learning and reinforcement learning (PPO) to this problem for the first time, comparing them against ILP. ILP achieves provably optimal solutions up to $19\times 19$ grids, while PatternBoost matches optimal performance up to $14\times 14$ grids with 96% test loss reduction. PPO achieves perfect solutions on $10\times 10$ grids but fails at $11\times 11$ grids, where constraint violations prevent valid configurations. These results demonstrate that classical optimization remains essential for exact solutions while AI methods offer competitive performance on smaller instances, with hybrid approaches presenting the most promising direction for scaling to larger problem sizes.
@@ 1 Introduction
A trivial upper bound of $2n$ points follows from the pigeonhole principle, as no row or column can contain more than two points. The first non-trivial lower bound was established by Paul Erdős, who used modular parabola constructions to place approximately $n$ points [ErdosRoth1951]. This was improved to approximately $1.5n$ by Hall, Jackson, Sudbery, and Wild using modular hyperbola constructions, which remains the best proven lower bound [Hall1975]. Exact solutions achieving $2n$ points have been found for all $n\leq 58$ [^1] T. Prellberg, private communication, 2025.. However, Guy and Kelly conjectured that for sufficiently large $n$, the maximum asymptotically approaches only $cn$ points, where $c\approx 1.814$ [GuyKelly1968]. This leaves a significant gap between the best-proven result and the conjectured limit, underscoring the problem’s persistent difficulty.
@@ 2.4.2 PPO Algorithm and Network Architecture
The policy network uses a multi-layer perceptron (MLP) with two hidden layers of 512 units each and ReLU activations. The architecture processes the flattened $n^{2}$-dimensional observation vector to produce action probabilities over the $n^{2}$ action space. The policy network (actor) $\pi_{\theta}(a|s)$ and value function (critic) $V_{\theta}(s)$ share feature extraction layers but have separate output heads, as shown in Figure 3.
@@ 3.2 Solution Quality and Scalability
Table 3 presents the solution quality achieved by each method across different grid sizes. We report the maximum number of points successfully placed while satisfying the no-three-in-line constraint (zero violations). For reference, we include known optimal values $2n$ for grid size $n\times n$.
@@
Proximal Policy Optimization. PPO achieved perfect solutions on $n=10$ grids (20 points with zero violations), successfully learning valid placement policies through reward-based feedback. On $n=11$ grids, the agent placed 22 points but incurred a single collinearity violation, indicating difficulty maintaining long-term constraint satisfaction as problem complexity increases. PPO demonstrated the ability to discover effective strategies from minimal problem structure (only the reward signal), but struggled to scale beyond small grid sizes where sequential planning becomes critical. Figures 7 and 7 illustrate this sharp performance transition between successful constraint satisfaction at $n=10$ and systematic failure at $n=11$, with detailed analysis of failure modes provided in Section 3.4.
@@ 3.4 Limitations
Figures 7 and 7 illustrate PPO’s performance transition between successful and failed constraint satisfaction. On $10\times 10$ grids (Figure 7), the agent consistently achieves perfect solutions with 20 points and zero violations, demonstrating successful policy learning for problems of this scale. However, on $11\times 11$ grids (Figure 7), despite placing 22 points, the agent incurs a single collinearity violation (highlighted point in red box), revealing the brittleness of learned policies as problem complexity increases. The violation occurs in a densely populated region where multiple collinearity constraints intersect, suggesting that the policy network struggles to maintain global constraint awareness when the number of placed points exceeds approximately 20. This represents a sharp transition from reliable performance to systematic failure as grid size increases by just one row and column.
@@ 4 Discussion and Conclusion
Future research should prioritize hybrid training paradigms that combine supervised pretraining (PatternBoost’s cross-entropy learning) with reinforcement fine-tuning (PPO’s exploratory optimization), mirroring successful approaches in language modeling. Transfer learning across grid sizes and curriculum strategies could enable training on smaller grids before scaling to larger instances, with meta-learning potentially discovering optimization strategies that generalize across problem scales. Classical-AI integration offers additional promising directions: ILP warm-starting with PatternBoost candidates could accelerate convergence on larger grids, while PatternBoost training on ILP-optimal smaller solutions could improve pattern quality beyond greedy baselines. Physics-inspired approaches present a particularly compelling direction, where combinatorial problems are reformulated as energy minimisation over relaxed Hamiltonians, enabling gradient-based neural network training [schuetz2022combinatorial].